OpenStudy (dannyo19):

What is the integral evaluated from 0 to (pi/4) of the integral evaluated from 0 to (pi/2) of sin(x+y)dxdy ?

OpenStudy (mathmale):

Hints: the inside integral is with respect to x, so x ranges from 0 to pi/2. Hold y constant (pretend that y is a constant). Find a way to integrate sin (x+y ) with respect to x.

OpenStudy (mathmale):

You may find it useful to use the trig identiy sin (a+b) = sin a cos b + cos a sin b, but it may also be possible to jump in right now and integrate sin (x + y) with respect to x.

OpenStudy (dannyo19):

So the integral of sinxcosy+cosxsiny then what?

OpenStudy (dannyo19):

arrh, the integral of sin(x+y) from zero to (pi/2) is sin(y)+cos(y)

OpenStudy (astrophysics):

\[\int\limits_{0}^{\pi/4}\int\limits_{0}^{\pi/2}\sin(x+y)dxdy\] this correct? Use the trig identity as mathmale mentioned\[\sin(a+b) = \sin(a)\cos(b)+\cos(a)\sin(b)\]

OpenStudy (astrophysics):

Yes, that sounds good

OpenStudy (astrophysics):

\[\int\limits_{0}^{\pi/4} siny+cosy dy\] this should be fairly simple

OpenStudy (astrophysics):